can you please tell me what does this mean: 17-( x3)=5
i have to figure out what =5 from taking 17 away from 3x something

Answer by hyh_124(6)

Dont get worried its pretty simple.
first analyse the situation
the question is 17-(x3)=5
multiply x3 by '-'
you will have -3x
now solve
17-3x=5 (shift 17 to the other side)
-3x =5-17
-3x =-12 (take minus common)
3x =12
x =12/3
x =4
Check
17-3x =5
17-3(4)=5
17-12 =5
5 =5